Can you help me with this :
y=3x+15
2x-3y= -17

Answer by ewatrrr(24785)


Hi
solving by substitution method:
y=3x+15
2x-3y= -17
2x - 3(3x+15) = -17
-7x = -17 + 45
-7x = 28
x = -4 and y = 3
CHECKING our Answer***
2*-4 - 3*3 = -8-9 = -17